# USTC-ICTS/PCFT-26-06

### v1

@@ Abstract
Recent observations from the DESI Collaboration indicate a preference for quintom dark energy, i.e., its equation of state evolves across the cosmological constant boundary w=1w=-1. It is well known that models with single perfect fluid or single scalar field minimally coupled to Einstein gravity develop perturbative instabilities around the crossing, thereby cannot realize the quintom scenario. In this paper, we provide a method to circumvent the instability problem of these models by introducing the coupling of dark energy to the Nieh-Yan density in teleparallel gravity. We show that with this coupling the background evolution is not affected, but the dark energy perturbation is removed from the menu of dynamical degrees of freedom, thus avoiding the intrinsic difficulties in the old models.
@@ I Introduction
Typical dynamical dark energy models, such as quintessence Ratra and Peebles (1988); Wetterich (1988); Caldwell et al. (1998) and phantom Caldwell (2002) models, cannot cross the cosmological constant boundary. The most important property of quintom dark energy Feng et al. (2005) is that its EoS crosses 1-1 during evolution. According to the direction of transition, the quintom dark energy can be classified as type A and type B models. For quintom-A, its EoS transited from the phase of w>1w>-1 to that of w<1w<-1 as universe expanding, while quintom-B has the opposite behavior. It is worth pointing out that current data, as shown in Abdul Karim and others (2025); Lodha and others (2025); Gu and others (2025) and other analyses, favor the quintom-B scenario. This challenges the dark energy model building. Simple models based on minimally coupled perfect fluid or kessence-like field Chiba et al. (2000); Armendariz-Picon et al. (2000) with single component cannot realize the quintom scenario, because around the time of w=1w=-1 crossing there emerges unavoidably unphysical instabilities in the dark energy perturbation. This has been noticed in Refs.Feng et al. (2005); Vikman (2005); Hu (2005); Zhao et al. (2005), and is formalized as a no-go theorem in Refs.Xia et al. (2008); Cai et al. (2010); Qiu (2010). As discussed in the next section, there are two kinds of instabilities involved in dark energy perturbation. One is the gradient instability caused by a negative sound speed square, another one is the ghost instability due to a kinetic term with wrong sign. The efforts to enable a w=1w=-1 crossing are made by introducing explicitly more degrees of freedom into the dark energy Feng et al. (2005); Hu (2005); Guo et al. (2005); Wei et al. (2005); Mughal and Ahmad (2021); Vázquez et al. (2024) or by considering higher derivatives in the scalar field model Li et al. (2005); Zhang and Qiu (2006); Cai et al. (2007). As shown in Ref.Li et al. (2012), quintom models without pathologies in perturbations can be constructed in terms of the scalar field with degenerate higher derivatives. In addition, dark energy models with EoS crossing 1-1 have been also discussed in some theories of modified gravity or non-minimal couplings Langlois et al. (2017, 2019); Horndeski (1974); Nicolis et al. (2009); Deffayet et al. (2009, 2011); Bamba et al. (2011); Cai et al. (2016); Arora and Sahoo (2022), and more recently in Refs.Nojiri et al. (2025a); Wang et al. (2025); Tsujikawa (2025); Chen et al. (2025); Wolf et al. (2025); Sánchez López et al. (2025); Toomey et al. (2025); Nojiri et al. (2025b); Tsujikawa (2026)
@@
In this paper, we propose a new method to circumvent the instability problem of single perfect fluid or single kessence-like field quintom models within the framework of teleparallel gravity Aldrovandi and Pereira (2013); Bahamonde et al. (2023). Teleparallel gravity provides a different picture for gravitation, in which the curvature and non-metricity tensor vanish and gravity is manifested by torsion. Model equivalent to general relativity (GR) can be constructed within this framework. In this paper we will extend this GR equivalent teleparallel gravity slightly by considering the coupling of dark energy to the Nieh–Yan density Nieh and Yan (1982). This formed the Nieh-Yan modified Teleparallel Gravity (NYTG) Li et al. (2020, 2021), which has been proposed as a consistent parity-violating gravity model. We will show that, since the Nieh-Yan density vanishes in the homogeneous and isotropic background of the universe, this coupling has no effect on the background evolution of dark energy, but it provides an additional constraint and forces the gauge-invariant dark energy perturbation to vanish. This provides a mechanism to remove the dark energy perturbation from the menu of dynamical degrees of freedom, and also to avoid the relevant difficulties in the old models. Consequently, one only needs to focus on the background evolution to construct quintom models with confidence.
@@
This paper is organized as follows. In Section II, we review the no-go behavior occurred in quintom models of single minimally coupled fluid or scalar field. To overcome these difficulties, we introduce the NYTG model in Section III, where dark energy is coupled to the Nieh-Yan density. The detailed studies on the fluid and scalar field dark energy models within NYTG framework are presented in sections IV and V, respectively. Our results demonstrate the absence of dynamical degree of freedom associated with dark energy perturbation. We list two toy models and illustrate their background evolutions, which exhibit the quintom-B behavior. Finally we conclude in Section VI.
@@ II Review on the no-go behavior of quintom model building
In this section, we review the no-go behavior occurred in quintom models of minimally coupled perfect fluid or scalar field with single component. We first consider the case of dark energy as a single perfect fluid, with minimal coupling to gravity, its action usually takes the form: Jackiw et al. (2004); De Felice et al. (2010); Schutz (1970); Schutz and Sorkin (1977):
@@
where $a(\eta)$ is the scale factor and $\eta$ is the conformal time. At this background there is only one independent equation for the fluid, i.e., the conservation law for the particle number, $n^{\prime}+3\mathcal{H}n=0$, where prime denotes the variation with respect to conformal time and $\mathcal{H}=a^{\prime}/a$. This means the number density scales as $n=n_{0}a^{-3}$ in the expanding universe and decreases monotonically with time. The continuity equation follows automatically,
$\rho^{\prime}+3\mathcal{H}(\rho+P)=f_{,n}\left(n^{\prime}+3\mathcal{H}n\right)=0,$ (7)
Since $\rho+P=\rho\,(1+w_{f})=nf_{,n}$, it is evident that $w_{f}>-1$ corresponds to $f_{,n}>0$, while $w_{f}<-1$ corresponds to $f_{,n}<0$. Therefore, the crossing point $w_{f}=-1$ occurs at a local extremum $f_{,n}=0$, and a sign change in $f_{,n}$ is required to realize quintom behavior. Accordingly, quintom-A is associated with $f_{,nn}>0$, while the quintom-B implies $f_{,nn}<0$. Since observations favor the quintom-B scenario, we only illustrate the evolution of the energy density $\rho(n)$ and the $P-\rho$ relation for the quintom-B scenario in FIG.2(a).
(a) The left panel illustrates the energy density $\rho(n)$ as a function of the particle number density $n$ in the quintom-B scenario. The right panel shows the $P-\rho$ relation in the same scenario. Arrows indicate the direction of evolution.
@@
As noted above, quintom behavior necessitates a sign change in $f_{{1,n}}$, which implies that $c_{s1}^{2}$ inevitably becomes negative once the EoS parameter of dark energy crosses w=1w=-1. Such a negative squared sound speed leads to unavoidable gradient instabilities, which in turn cause the system to be physically unstable. Furthermore, as mentioned before, the quintom-B scenario requires $f_{1{,nn}}<0$, leading to a negative $z_{1}^{2}$. This means the dark energy perturbation $\zeta_{1}$ is a ghost mode because it has a kinetic term with wrong sign. The existence of ghost mode makes the Hamiltonian unbounded from below, leading to severe problems in quantization.
@@
Additionally, the possibility of realizing a quintom scenario must be excluded when the Lagrangian density depends only on $X$: $\mathcal{L}=\mathcal{L}(X)$. In this case, the model possesses the shift symmetry, $\phi\rightarrow\phi+{\rm constant}$, exactly. The corresponding conservation law for the Noether charge is $\mathring{\nabla}_{\mu}j^{\mu}=0$ with $j^{\mu}=P_{,X}\nabla^{\mu}\phi=\frac{\rho+P}{\sqrt{2X}}U^{\mu}$. One can read from this current the charge density $n=\frac{\rho+P}{\sqrt{2X}}$, which scales as $n\propto a^{-3}$ in the expanding universe, just the same as the particle number density in the fluid case. So that $n\propto\rho+P$ cannot cross zero point, i.e., $w$ cannot cross the boundary 1-1. Therefore, to have a quintom model the scalar field $\phi$ must appear explicitly in the Lagrangian, i.e., $\mathcal{L}=\mathcal{L}(X,\phi)$.
@@ III The dark energy in NYTG
The review in the previous section told us that, in Einstein gravity with minimal couplings, the unavoidable instabilities in perturbations prevent either a single perfect fluid or a single kessence-like scalar field from realizing the quintom scenario. Now we propose a possible solution to this problem within the framework of teleparallel gravity. In our method, the quintom dark energy is still represented by a single perfect fluid or a single kessence-like scalar field. The key point is that we will introduce a coupling of the dark energy to the Nieh-Yan density. In this way, we incorporated the dark energy into the NYTG model, and as will be shown below, the original instabilities in dark energy perturbations are eliminated in this model.
@@
As previously stated, our model extend GR equivalent teleparallel gravity slightly by considering the coupling of dark energy to the Nieh–Yan density defined as $T_{\lambda\mu\nu}\tilde{T}^{\lambda\mu\nu}=(1/2)\varepsilon^{\mu\nu\rho\sigma}T^{\lambda}_{\phantom{1}\mu\nu}T_{\lambda\rho\sigma}\,$. Including dark energy and other matter components, the action under consideration is:
@@ IV.2.2 Quadratic action for scalar perturbations
The subsequent step is to realize the quintom-B behavior using an appropriate single perfect fluid model. In illustrating the evolution of the EoS parameter $w_{{}_{DE}}$, we have assumed $\Omega_{m,0}\approx 0.31$ and $h\approx 0.69$. The background evolution can be rewritten as follows:
@@
where the dot denotes a derivative with respect to physical time $t$ and $H=\dot{a}/a$. The final equation indicates that the number density $n_{1}$ scales as $n_{1}=n_{0}a^{-3}$, and can therefore be expressed in terms of $a$. Accordingly, $f_{1}$ is also a positive function of $a$. In order to realize the quintom-B scenario, the function $f_{1}$ is chosen to have the following form:
$\rho_{{}_{DE}}=f_{1}(a)=f_{0}\left(\frac{a}{a_{i}}\right)^{-3(w_{0}+w_{a}+1)}e^{3w_{a}(a-a_{i})},$ (44)
where $a_{i}\approx 0.001$ is the initial value of $a$, $w_{0}$ and $w_{a}$ are non-zero constant. Thus the EoS parameter $w_{{}_{DE}}$ is
$w_{{}_{DE}}=w_{0}+w_{a}(1-a).$ (45)
It can be shown that the parameters $w_{0}$ and $w_{a}$ have the capacity to affect the evolution of the EoS parameter $w_{{}_{DE}}$, while $f_{0}$ has an influence on the magnitude of $\rho_{{}_{DE}}$. The evolution of $w_{{}_{DE}}$ as a function of redshift $z$ is presented in FIG.2, where we adopt $w_{0}=-0.67,\,w_{a}=-1.09$ and $f_{0}/M^{2}=3.96\times 10^{-6}$. Here $M\approx 8.77\times 10^{-61}$ is a rescaling parameter introduced for computational convenience. As can be seen, $w_{{}_{DE}}(z)$ evolves from values below 1-1 to above 1-1 and crossing the cosmological constant boundary w=1w=-1 at $z\approx 0.44$. At the present epoch, i.e., $z=0$, the model predicts $w_{0}=-0.67$.
Figure 2: The evolution of $w_{{}_{DE}}(z)$ of single fluid dark energy with $w_{0}=-0.67,\,w_{a}=-1.09$ and $f_{0}/M^{2}=3.96\times 10^{-6}$. The gray dashed line marks the cosmological constant $\Lambda$ with EoS parameter w=1w=-1.
(a) The way $w_{{}_{DE}}(z)$ changes with different values of a certain parameter. Left panel corresponds to different values of $w_{0}$ and right panel is associated with different values of $w_{a}$. The gray dashed line in both of panels marks the cosmological constant $\Lambda$ with EoS parameter w=1w=-1.
The effects of $w_{0}$ and $w_{a}$ on the evolution of $w_{{}_{DE}}(z)$ are presented in FIG.4(a). The left panel shows the behavior of $w_{{}_{DE}}(z)$ for a range of $w_{0}$ values with $w_{a}$ fixed at $w_{a}=-1.09$. It is evident that lower values of $w_{0}$ shift the w=1w=-1 crossing to smaller redshift and reduce the present value $w_{0}$. With $w_{0}$ fixed at $w_{0}=-0.67$, the right panel illustrates the effect of parameter $w_{a}$ on $w_{{}_{DE}}(z)$. Reducing $w_{a}$ shifts the crossing toward lower $z$, while the value of $w_{0}$ remains unchanged.

### v2

@@ Abstract
Recent observations from the DESI Collaboration indicate a preference for quintom dark energy, i.e., its equation of state evolves across the cosmological constant boundary w=1w=-1. It is well known that models with single perfect fluid or single scalar field minimally coupled to Einstein gravity develop perturbative instabilities around the crossing, thereby cannot realize the quintom scenario. In this paper, we propose a method to circumvent the instability problem of these models by considering the coupling of dark energy to the Nieh-Yan density within the framework of teleparallel gravity. We show that with this coupling the background evolution is not affected, but the dark energy perturbation is removed from the menu of dynamical degrees of freedom, thus avoiding the inherent difficulties in the old models. Furthermore, the Nieh-Yan coupling causes parity violation in gravitational waves, and this can be considered as a clear prediction of this mechanism.
@@ I Introduction
Typical dynamical dark energy models, such as quintessence Ratra and Peebles (1988); Wetterich (1988); Caldwell et al. (1998) and phantom Caldwell (2002) models, cannot cross the cosmological constant boundary. The most important property of quintom dark energy Feng et al. (2005) is that its EoS crosses 1-1 during evolution. According to the direction of transition, the quintom dark energy can be classified as type A and type B models. For quintom-A, its EoS transited from the phase of w>1w>-1 to that of w<1w<-1 as universe expanding, while quintom-B has the opposite behavior. It is worth pointing out that current data, as shown in Abdul Karim and others (2025); Lodha and others (2025); Gu and others (2025) and other analyses, favor the quintom-B scenario. This challenges the dark energy model building. Simple models based on minimally coupled perfect fluid or kessence-like field Chiba et al. (2000); Armendariz-Picon et al. (2000) with single component cannot realize the quintom scenario, because around the time of w=1w=-1 crossing there emerges unavoidably unphysical instabilities in the dark energy perturbation. This has been noticed in Refs.Feng et al. (2005); Vikman (2005); Hu (2005); Zhao et al. (2005), and is formalized as a no-go theorem in Refs.Xia et al. (2008); Cai et al. (2010); Qiu (2010). As discussed in the next section, there are two kinds of instabilities involved in dark energy perturbation. One is the gradient instability caused by a negative sound speed square, another one is the ghost instability due to a kinetic term with wrong sign. The efforts to enable a w=1w=-1 crossing are made by introducing explicitly more degrees of freedom into the dark energy Feng et al. (2005); Hu (2005); Guo et al. (2005); Wei et al. (2005); Mughal and Ahmad (2021); Vázquez et al. (2024) or by considering higher derivatives in the scalar field model Li et al. (2005); Zhang and Qiu (2006); Cai et al. (2007). As shown in Ref.Li et al. (2012), quintom models without pathologies in perturbations can be constructed in terms of the scalar field with degenerate higher derivatives. In addition, dark energy models with EoS crossing 1-1 have been also discussed in some theories of modified gravity or non-minimal couplings Langlois et al. (2017, 2019); Horndeski (1974); Nicolis et al. (2009); Deffayet et al. (2009, 2011); Bamba et al. (2011); Cai et al. (2016); Arora and Sahoo (2022), and more recently in Refs.Basilakos et al. (2025); Paliathanasis (2025); Li et al. (2026a); Nojiri et al. (2025a); Wang et al. (2025); Tsujikawa (2025); Chen et al. (2025); Wolf et al. (2025); Sánchez López et al. (2025); Toomey et al. (2025); Paliathanasis et al. (2025); Nojiri et al. (2025b); Li et al. (2026b); Tsujikawa (2026)
@@
In this paper we address the theoretical challenge of realizing the quintom scenario with single perfect fluid or single kessence-like field. We propose a new method to circumvent the instability of these quintom models within the framework of teleparallel gravity Aldrovandi and Pereira (2013); Bahamonde et al. (2023). Teleparallel gravity provides a different picture for gravitation, in which the curvature and non-metricity tensor vanish and gravity is manifested by torsion. Model equivalent to general relativity (GR) can be constructed within this framework. In this paper we will extend this GR equivalent teleparallel gravity slightly by considering the coupling of dark energy to the Nieh–Yan density Nieh and Yan (1982). This formed the Nieh-Yan modified Teleparallel Gravity (NYTG) Li et al. (2020, 2021), which has been proposed as a consistent and healthy parity-violating gravity model. We will show that, since the Nieh-Yan density vanishes in the homogeneous and isotropic background of the universe, this coupling has no effect on the background evolution of dark energy, but it provides an additional constraint and forces the gauge-invariant dark energy perturbation to vanish. This provides a mechanism to remove the dark energy perturbation from the menu of dynamical degrees of freedom, and also to avoid the inherent difficulties in the old models. Consequently, one only needs to focus on the background evolution to construct quintom models with confidence. In addition, the incorporation of dark energy into the NYTG model inevitably leads to parity violation in gravitational waves, and this can be considered as unique theoretical prediction of this mechanism, further signals the dynamical nature of dark energy.
@@
This paper is organized as follows. In Section II, we review the no-go behavior occurred in quintom models of single minimally coupled fluid or kessence-like scalar field. To overcome these difficulties, we introduce the NYTG model in Section III, where dark energy is coupled to the Nieh-Yan density. The detailed studies on the fluid and scalar field dark energy models within NYTG framework are presented in sections IV and V, respectively. Our results demonstrate the absence of dynamical degree of freedom associated with dark energy perturbation. We list two toy models and illustrate their background evolutions, which exhibit the quintom-B behavior. Section VI presents an additional demonstration that our mechanism also predicts parity violation in gravitational waves. Finally we conclude in Section VII.
@@ II Review on the no-go behavior of quintom model building
In this section, we review the no-go behavior occurred in quintom models of minimally coupled perfect fluid or scalar field with single component. We first consider the case of dark energy as a single perfect fluid, with minimal coupling to Einstein gravity, its action usually takes the form: Jackiw et al. (2004); De Felice et al. (2010); Schutz (1970); Schutz and Sorkin (1977):
@@
where $a(\eta)$ is the scale factor and $\eta$ is the conformal time. At this background there is only one independent equation for the fluid, i.e., the conservation law for the particle number, $n^{\prime}+3\mathcal{H}n=0$, where prime denotes the variation with respect to conformal time and $\mathcal{H}=a^{\prime}/a$. This means the number density scales as $n=n_{0}a^{-3}$ in the expanding universe and decreases monotonically with time, here the subscript 0 denotes the present value and we have also used the normalization $a_{0}=1$. The continuity equation follows automatically,
$\rho^{\prime}+3\mathcal{H}(\rho+P)=f_{,n}\left(n^{\prime}+3\mathcal{H}n\right)=0.$ (7)
One of the advantages of the perfect fluid model is one can know the specific form of the theoretical model from the observational data strait forwardly. Once the dependence of EoS on the scale factor $w_{f}(a)$ is reconstructed from data, we can obtain the function $\rho=f(n)$ through the integration (5) and the relation $a=(n_{0}/n)^{1/3}$.
Now we turn to the quintom behavior of the perfect fluid model. Since $\rho+P=\rho\,(1+w_{f})=nf_{,n}$, it is evident that $w_{f}>-1$ corresponds to $f_{,n}>0$, while $w_{f}<-1$ corresponds to $f_{,n}<0$. Therefore, the crossing point $w_{f}=-1$ occurs at a local extremum $f_{,n}=0$, and a sign change in $f_{,n}$ is required to realize quintom behavior. Accordingly, quintom-A is associated with $f_{,nn}>0$, while the quintom-B implies $f_{,nn}<0$. The evolution of the energy density $\rho(n)$ and the $P-\rho$ relations for the quintom-A and quintom-B scenarios are illustrated in FIG.2(a) and FIG.2(b), respectively.
(a) The left panel illustrates the energy density $\rho(n)$ as a function of the particle number density $n$ in the quintom-A scenario. The right panel shows the $P-\rho$ relation in the same scenario. Arrows indicate the direction of evolution.
@@
(b) The left panel illustrates the energy density $\rho(n)$ as a function of the particle number density $n$ in the quintom-B scenario. The right panel shows the $P-\rho$ relation in the same scenario. Arrows indicate the direction of evolution.
@@
As noted above, the quintom behavior necessitates a sign change in $f_{{1,n}}$, which implies that $c_{s1}^{2}$ inevitably becomes negative once the EoS parameter of dark energy crosses w=1w=-1. Such a negative squared sound speed leads to unavoidable gradient instabilities, which in turn cause the system to be physically unstable. Furthermore, as mentioned before, the quintom-B scenario requires $f_{1{,nn}}<0$, leading to a negative $z_{1}^{2}$. This means the dark energy perturbation $\zeta_{1}$ is a ghost mode because it has a kinetic term with wrong sign. The existence of ghost mode makes the Hamiltonian unbounded from below, leading to severe problems in quantization.
@@
Additionally, the possibility of realizing a quintom scenario must be excluded when the Lagrangian density depends only on $X$: $\mathcal{L}=\mathcal{L}(X)$. In this case, the model possesses an exact shift symmetry, $\phi\rightarrow\phi+{\rm constant}$. The corresponding conservation law for the Noether charge is $\mathring{\nabla}_{\mu}j^{\mu}=0$ with $j^{\mu}=P_{,X}\nabla^{\mu}\phi=[(\rho+P)/\sqrt{2X}]U^{\mu}$. One can read from this current the charge density $n=(\rho+P)/\sqrt{2X}$, which scales as $n\propto a^{-3}$ in the expanding universe, just the same as the particle number density in the fluid case. So that $n\propto\rho+P$ cannot cross the zero point, this means $w$ cannot cross the boundary 1-1. Therefore, to have a quintom model the scalar field $\phi$ must appear explicitly in the Lagrangian, i.e., $\mathcal{L}=\mathcal{L}(X,\phi)$.
@@ III The dark energy in NYTG
The review in the previous section told us that, in Einstein gravity with minimal couplings, the unavoidable instabilities in perturbations prevent either a single perfect fluid or a single kessence-like scalar field from realizing the quintom scenario. In order to address this theoretical problem, we propose a solution within the framework of teleparallel gravity. In our method, the quintom dark energy is still represented by a single perfect fluid or a single kessence-like scalar field. The key point is that we will introduce a coupling of the dark energy to the Nieh-Yan density. In this way, we incorporated the dark energy into the NYTG model, and as will be shown below, the original instabilities in dark energy perturbations are eliminated in this model.
@@
As previously mentioned, our model extends GR equivalent teleparallel gravity slightly by considering the coupling of dark energy to the Nieh–Yan density defined as $T_{\lambda\mu\nu}\tilde{T}^{\lambda\mu\nu}=(1/2)\varepsilon^{\mu\nu\rho\sigma}T^{\lambda}_{\phantom{1}\mu\nu}T_{\lambda\rho\sigma}\,$. Including dark energy and other matter components, the action under consideration is:
@@
Similar to the Chern-Simons term studied frequently in the literature, the Nieh-Yan density is also topological density and appears as a total derivative of the tetrad. The coupling $\phi\,T_{\lambda\mu\nu}\tilde{T}^{\lambda\mu\nu}$ is invariant under the shift of the field $\phi\rightarrow\phi+\phi_{0}$ by a constant $\phi_{0}$, up to a surface term. Such kind of couplings are familiar for the axion-like fields. The axion-like fields are pseudo-Goldstone bosons and can only have the couplings preserving the shift symmetry. In this sense, we may treat the dark energy field $\phi$ as axion-like, so that its coupling to the Nieh-Yan density is highly expected within the framework of teleparallel gravity, very similar to the interactions of QCD axion with gauge fields.
@@ IV.2.2 Quadratic action for scalar perturbations
The subsequent step is to realize the quintom-B behavior using an appropriate single perfect fluid model. We have assumed $\Omega_{m,0}\approx 0.31$ and $h\approx 0.69$. The background evolution can be rewritten as follows:
@@
where the dot denotes a derivative with respect to physical time $t$ and $H=\dot{a}/a$. For simplicity, we consider the toy model in which the EoS of dark energy depends on the scale factor exactly follow the Chevallier-Polarski-Linder (CPL) parametrization
$w_{{}_{DE}}=w_{0}+w_{a}(1-a),$ (44)
where $w_{0}$ and $w_{a}$ are two constants determined by observational data. As mentioned before, we can obtain the model easily:
$\rho_{{}_{DE}}=f_{1}(n_{1})=f_{0}\left(\frac{n_{1}}{n_{0}}\right)^{w_{0}+w_{a}+1}e^{3w_{a}[(n_{0}/n_{1})^{1/3}-1]},$ (45)
where $n_{1}=n_{0}a^{-3}$ and $f_{0}$ is the present dark energy density.
The dependence of $f_{1}(n_{1})/f_{0}$ on $\mathrm{ln}(n_{1}/n_{0})$ and the evolution of $w_{{}_{DE}}$ as a function of redshift $z$ are presented in FIG.2, where we adopt $w_{0}=-0.667,\,w_{a}=-1.09$ and $f_{0}\approx 8.63\times 10^{-61}$. The values of $w_{0}$ and $w_{a}$ are taken from the best-fit results of the observational constraints on the CPL parametrization, as derived in Ref.Abdul Karim and others (2025) using the combined “DESI+CMB+Union3” data. As shown in FIG.2, the function $f_{1}(n_{1})/f_{0}$ peaks at $n_{1}/n_{0}\approx 2.99$. As a comparison, the present value $f_{1}(n_{1})/f_{0}=1$ at $n_{1}/n_{0}=1$ is also indicated in the figure. Since the peak is associated with the crossing point, one can evaluate that the dark energy EoS $w_{{}_{DE}}$ crosses the cosmological constant boundary w=1w=-1 at the redshift $z\approx 0.44$. In addition, this toy model predicts that the dark energy density will decay to zero in the far future.
Figure 2: Left panel: The evolution of $f_{1}(n_{1})/f_{0}$ of single fluid dark energy with $w_{0}=-0.667,\,w_{a}=-1.09$. The peak of $f_{1}(n_{1})/f_{0}$ corresponds to the crossing point, it is very close to the present time. Right panel: The evolution of $w_{{}_{DE}}(z)$ of single fluid dark energy. The gray dashed line marks the cosmological constant $\Lambda$ with EoS parameter w=1w=-1.